# Bayesian Transfer Operators in Reproducing Kernel Hilbert Spaces

### v1

@@ 2.2 Sparse Variational Bayesian Transfer Operators
Compressing the exact posterior to a smaller basis dictionary $\Phi_{Z}$ is possibly the easiest step, and demonstrates how straightforward it is to apply Gaussian process techniques to kernel transfer operators. Because, once we have framed the problem within a Bayesian framework, we can simply apply the variational free energy (VFE) method (see Appendix A.3 and Titsias (2009)):
@@
where $\tilde{\mathbf{C}}_{XX}\!\coloneqq\!\mathbf{K}_{ZX}\mathbf{K}_{ZX}^{\top}\!+\!\sigma^{2}_{Y}\mathbf{K}_{ZZ}$ is the noisy Gramian.
In addition to providing a means of learning a sparse dictionary, the VFE method can be used for hyperparameter optimization. Hyperparameter tuning is another obstacle for kernel-based Koopman methods; one should ideally design the reproducing kernel function such that the Koopman operator is represented in an RKHS that is both densely defined and closable (Ikeda et al., 2022). Consequently, finding such a kernel function tends to be a delicate and time-consuming task.
@@ 2.3 Koopman Mode Decomposition
Notice that since $\mathbf{C}_{XY}\coloneqq\mathbf{K}_{ZX}\mathbf{K}_{ZY}$, (2.5) recovers the standard algorithm for EDMD (Klus et al., 2016). In other words, the Bayesian approach we have followed produces the same expression as the least-squares approach of EDMD with a Tikhonov parameter, $\sigma_{Y}^{2}$. Of course, since EDMD is valid for more general choices of dictionaries, the previous statement is only true for a kernel-based dictionary (Williams et al., 2015a).
@@
where $\mathbf{\Lambda}\in\mathbb{C}^{M\times M}$ is a diagonal matrix containing the discrete-time eigenvalues. The right eigenvectors $\mathbf{W}\in\mathbb{C}^{M\times M}$ contain the projection vectors onto the eigenfunction basis, whereas the left eigenvectors ${\mathbf{V}}_{\kappa}=\mathbf{W}^{-1}$ correspond to the modes of the observables $\mathbf{k}_{Z}:\mathbb{X}\rightarrow\mathbb{R}^{M}$.
@@
The implication of (2.7) is that for a single-step prediction, the model in (2.10) is also the posterior mean of a GP prior placed on the flow map. The point is that we can perform DMD on a random variable associated with the Koopman operator.
@@ 2.4 Prediction Theory
$\mathbf{k}_{Z}(X_{k})=\mathbf{U}^{\top}\mathbf{k}_{Z}(X_{k-1})\,+\,{\mathbf{L}}_{\textrm{pst}}(X_{k-1})\bm{\omega}_{k},$ (2.8)
where $\bm{\omega}_{k}\sim\mathcal{N}^{M}(\bm{0},\mathbf{I}_{M})$. The process $\{\bm{\omega}_{i}\}_{i=1}^{k}$ is an i.i.d. martingale difference sequence adapted to the sequence of information sets $\{J_{i}\}_{i=0}^{k}$. Each information set $J_{k}$ represents all the information available to an agent at time-step $k$, and contains all measurable functions of $\{X_{0},\bm{\omega}_{1},\ldots,\bm{\omega}_{k}\}$.
The lower triangular matrix ${\mathbf{L}}_{\textrm{pst}}(X_{k})$ is determined by the Cholesky decomposition of ${\mathbf{\Xi}}_{\textrm{pst}}^{1}(X_{k})$. This covariance matrix encodes the one-step-ahead forecasting error of $\mathbf{k}_{Z}(X_{k})$ conditioned on $J_{k-1}$. In the current setting this matrix is ${\mathbf{\Xi}}_{\textrm{pst}}^{1}(X_{k})={\kappa}_{\textrm{pst}}(X_{k},X_{k})\,\mathbf{K}_{\textrm{bc}}$. Now recall that the expression for the sparse posterior kernel (see (A.6) in Appendix A.3) is
@@
where $\tilde{\mathbf{B}}\coloneqq\left(\mathbf{K}_{ZZ}^{-1}-\sigma_{Y}^{2}\tilde{\mathbf{C}}_{XX}^{-1}\right)$. Hence, the single-step prediction error is represented by the discretized version of the posterior kernel $\xi_{\textrm{pst}}(\cdot,\cdot)$ and subsequently the operator $\mathcal{C}_{\textrm{bc}}$.
@@
$\begin{split}\mathbf{k}_{Z}(X_{k})&=\left(\mathbf{U}^{k}\right)^{\top}\mathbf{k}_{Z}(X_{0})+\sum_{j=0}^{k-1}\left(\mathbf{U}^{j}\right)^{\top}{\mathbf{L}}_{\textrm{pst}}(X_{k-j-1})\bm{\omega}_{k-j}.\end{split}$
@@
$\begin{split}\hat{\bm{x}}_{k}&=\mathbb{E}\left[X_{k}\mid X_{0}=\bm{x}_{0}\right]=\mathbf{A}_{k}^{\top}\,\mathbf{k}_{Z}(\bm{x}_{0}),\end{split}$ (2.10)
where the weight matrix is $\mathbf{A}_{k}\coloneqq\mathbf{W}\,\mathbf{\Lambda}^{k}\,{\mathbf{V}}_{f}\in\mathbb{R}^{M\times D}$. In deriving (2.10) we used the relationship $X_{k}=(\mathbf{W}{\mathbf{V}}_{f})^{\top}\mathbf{k}_{Z}(X_{k})$, and the fact that a linear transformation of a Gaussian random vector yields another Gaussian random vector (Peebles Jr, 2001). This property was also the key to calculating the Bayesian-consistency matrix $\mathbf{K}_{\textrm{bc}}\in\mathbb{R}^{M\times M}$.
@@
$\begin{split}\mathbf{K}_{\textrm{pst}}^{k}(\bm{x})+\sigma_{Y}^{2}\mathbf{I}_{D}&=\mathbf{A}_{0}^{\top}\left({\mathbf{\Xi}}_{\textrm{pst}}^{k}(\bm{x})+\sigma_{Y}^{2}\mathbf{K}_{\textrm{bc}}\right)\mathbf{A}_{0}.\end{split}$ (2.11)
Given the independence assumption of the components of the flow map, the single-step covariance matrix with respect to the full-state observables is $\mathbf{K}_{\textrm{pst}}^{1}=\kappa_{\textrm{pst}}(\bm{x},\bm{x})\,\mathbf{I}_{D}$. Hence, it follows that discretizing the operator $\mathcal{C}_{\textrm{bc}}$ leads to
$\begin{split}\mathbf{K}_{\textrm{bc}}&=\left(\mathbf{A}_{0}\mathbf{A}_{0}^{\top}\right)^{\dagger}={\mathbf{V}}_{\kappa}^{*}\left(\mathbf{V}_{f}\mathbf{V}_{f}^{*}\right)^{\dagger}{\mathbf{V}}_{\kappa}.\end{split}$
Since the matrix $\mathbf{K}_{\textrm{bc}}$ always has a rank less than or equal to $D$, we employed the Moore–Penrose pseudoinverse . Fortunately, well-posedness is already ensured by (2.7).
@@
$\zeta_{k}\coloneqq\mathbf{k}_{Z}(X_{k})-\mathbf{k}_{Z}(\hat{\bm{x}}_{k})=\sum_{j=0}^{k-1}\left(\mathbf{U}^{j}\right)^{\top}{\mathbf{L}}_{\textrm{pst}}(X_{k-j-1})\bm{\omega}_{k-j},$
@@
$=\mathbb{E}\left[\sum_{j=0}^{k-1}\left(\mathbf{U}^{j}\right)^{\top}\;\mathbf{\Xi}_{\textrm{pst}}^{1}\left(X_{k-j-1}\right)\;\mathbf{U}^{j}\mid J_{k}\right].$
As this expression is in general not analytically tractable, we suggest approximating it by performing a second-order Taylor series expansion about the predicted mean:
@@
which we can interpret as iteratively accumulating the prediction errors arising from random ‘shocks’ at each intermediate step. The magnitude of these shocks are dependent on the state of the system at each step, hence the heteroskedasticity.
@@ 2.5 Model Selection
Hyperparameter optimization proceeds as if one were only interested in learning the optimal feature space for the flow map with the VFE method (Titsias, 2009). This is an optimization problem that is solved by maximizing the sum of the VFEs of each of the components of the flow map $\bm{F}=\{f_{i}\}_{i=1}^{D}$. The reasoning is that the theoretical Bayesian structure of our model also ensures a coherent parameterization of the Koopman and Perron–Frobenius operators.
@@
Thereby, we used the approximate linear dependence (ALD) selection criterion (Engel et al., 2004) as a first pass to determine a subset of candidate locations for the pseudo-inputs from the training inputs. Since the ALD approach is a local ‘whack-a-mole’ active learning scheme, we followed it up by implementing active learning Cohn (Seo et al., 2000): a more globally oriented methodology that approximates a full A-optimal design (Gramacy, 2020). As a last step, only after a predefined number of new dictionary functions had been identified was the VFE hyperparameter tuning algorithm activated, with the greedy results given as a warm start. This process is interleaved with the adaptation of the generative hyperparameters, and repeats until no more pseudo-inputs are added to the basis dictionary or a threshold is passed, e.g., a limit on the dictionary size.
@@ 3 Numerical Experiments
With the conceptual unification of GP regression and DMD complete, we present the results obtained from a few experiments. In all of the experiments, we used the $\text{Matérn}^{5/2}$ kernel (Matérn, 2013) with automatic relevance determination (ARD; Rasmussen and Williams (2006)):
@@
For all numerical experiments, the data was standardized, and we emphasize that the data collection setup was asymmetric with respect to measurement noise. Since our model is not theoretically equipped to account for sensor noise on the inputs, we restricted the corruption to the dynamically evolved states. This simplification is admissible in surrogate modeling, but unrealistic in physical data acquisition contexts.
@@ 3.1 Generalization Errors
We compared the multi-step errors of our algorithm (which we referred to as GP-TCCA in Appendix D) using the Van der Pol oscillator as a benchmark:
@@
where $\alpha=2$ controls the degree of . The dynamical system was simulated by a numerical ODE solver and sampled in time intervals of $\Delta t=50.0\text{ [ms]}$. Figure 1 depicts an example trajectory and its associated forecast.
@@
[Uncaptioned image][Uncaptioned image]
@@
We also compared the VFE method to the ‘variational approach to Markov processes’ (VAMP) for hyperparameter optimization (Wu and Noé, 2020). Specifically we maximized the VAMP-$2$ score, i.e., the total kinetic variance of the model (Noé and Clementi, 2015).
When optimizing with the VAMP score, the two DMD algorithms are virtually indistinguishable, suggesting that the Tikhonov regularization in our model has little to no effect. Although this added robustness is desirable, we also observe that maximizing the variational free energy of the flow map yields higher short-term accuracies. This suggests that our model could be improved by explicitly enforcing more robust dynamical features and by augmenting the objective with information from the inverse flow map, i.e., adopt a forward–backward perspective on the dynamics (Azencot et al., 2020; Klus et al., 2019).
Figure 2 further shows that our DMD variant outperforms Exact EDMD under noisy conditions and when the VFE objective is maximized. This aligns with our expectations given the explicit treatment of sensor noise and the additional structure imposed by the Bayesian model. However, the nonlinear growth did prompt a closer examination of the regularization strategy in our GP-DMD model.
[Uncaptioned image][Uncaptioned image]

### v2

@@ 2.2 Sparse Variational Bayesian Transfer Operators
Compressing the exact posterior to a smaller basis dictionary $\Phi_{Z}\coloneqq\{\varphi\left(\bm{z}_{i}\right)\}_{i=1}^{M}\in\mathbb{H}^{1\times M}$ with $M<N$, is possibly the easiest step, and demonstrates how straightforward it is to apply Gaussian process techniques to kernel transfer operators. Because, once we have framed the problem within a Bayesian framework, we can simply apply the variational free energy (VFE) method (see Appendix A.3 and Titsias (2009)):
@@
where we have the compressed covariance operators $\Xi_{ZX}\,\coloneqq\,\mathbf{K}_{ZX}\,\otimes\,\mathcal{C}_{\textrm{bc}}$ and $\Xi_{ZZ}\,\coloneqq\,\mathbf{K}_{ZZ}\,\otimes\,\mathcal{C}_{\textrm{bc}}$. The sparse covariance matrices $\mathbf{K}_{ZX}\,\coloneqq\,\Phi_{Z}^{\top}\,\Phi_{X},\;\mathbf{K}_{ZZ}\,\coloneqq\,\Phi_{Z}^{\top}\,\Phi_{Z}$ allow us to define the noisy Gramian $\tilde{\mathbf{C}}_{XX}\,\coloneqq\,\mathbf{K}_{ZX}\,\mathbf{K}_{ZX}^{\top}\,+\,\sigma^{2}_{Y}\mathbf{K}_{ZZ}$. Lastly the feature vectors are $\mathbf{k}_{Z}(\bm{x})\!\coloneqq\!\Phi_{Z}^{\top}\varphi(\bm{x})$ and $\xi_{Z}({\bm{x}})\!\coloneqq\!\mathbf{k}_{Z}(\bm{x})\otimes\mathcal{C}_{\textrm{bc}}$.
In addition to providing a means of learning a sparse dictionary, the VFE method offers convergence-rate guarantees (Burt et al., 2019) and can also be used for hyperparameter optimization. Hyperparameter tuning is another obstacle for kernel-based Koopman methods, since one should ideally design the reproducing kernel function such that the Koopman operator is represented in an RKHS that is both densely defined and closable (Ikeda et al., 2022). Consequently, finding such a kernel function tends to be a delicate and time-consuming task.
@@ 2.3 Koopman Mode Decomposition
Notice that since $\mathbf{C}_{XY}\coloneqq\mathbf{K}_{ZX}\mathbf{K}_{ZY}^{\top}$ and $\mathbf{K}_{ZY}\coloneqq\Phi_{Z}^{\top}\,\Phi_{Y}$, equation (2.5) recovers the standard algorithm for EDMD (Klus et al., 2016). In other words, the Bayesian approach we have followed produces the same expression as the least-squares approach of EDMD with a Tikhonov parameter, $\sigma_{Y}^{2}$. Of course, since EDMD is valid for more general choices of dictionaries, the previous statement is only true for a kernel-based dictionary (Williams et al., 2015a).
@@
where $\mathbf{\Lambda}\in\mathbb{C}^{M\times M}$ is a diagonal matrix containing the discrete-time eigenvalues. The right eigenvectors $\mathbf{W}\in\mathbb{C}^{M\times M}$ contain the projection vectors onto the eigenfunction basis, whereas the left eigenvectors ${\mathbf{V}}_{\kappa}=\mathbf{W}^{-1}$ correspond to the modes of the observables $\mathbf{k}_{Z}(\cdot)=[\kappa_{\textrm{pr}}(\bm{z}_{1},\cdot),\,\ldots,\,\kappa_{\textrm{pr}}(\bm{z}_{M},\cdot)]^{\top}$.
@@
The implication of (2.7) is that for a single-step prediction, the model in (2.10) is also the posterior mean of a GP prior placed on the flow map (using the dynamically evolved states $\mathbf{Y}$ as the targets). The point is that we can perform DMD on a random variable associated with the Koopman operator while maintaining consistency with the nonlinear dynamics on the state manifold.
@@ 2.4 Prediction Theory
$\mathbf{k}_{Z}(X_{k})=\mathbf{A}_{\kappa}^{1}\;\mathbf{k}_{Z}(X_{k-1})\,+\,{\mathbf{L}}_{\textrm{pst}}(X_{k-1})\bm{\omega}_{k},$ (2.8)
where $\bm{\omega}_{k}\sim\mathcal{N}^{M}(\bm{0},\mathbf{I}_{M})$ and the multi-step transition matrix is $\mathbf{A}_{\kappa}^{k}\coloneqq\left(\mathbf{W}\,\mathbf{\Lambda}^{k}\,{\mathbf{V}}_{\kappa}\right)^{\top}$. The process $\{\bm{\omega}_{i}\}_{i=1}^{k}$ is an i.i.d. martingale difference sequence adapted to the sequence of information sets $\{J_{i}\}_{i=0}^{k}$. Each information set $J_{k}$ represents all the information available to an agent at time-step $k$, and contains all measurable functions of $\{X_{0},\bm{\omega}_{1},\ldots,\bm{\omega}_{k}\}$.
The lower triangular matrix ${\mathbf{L}}_{\textrm{pst}}(X_{k})$ is determined by the Cholesky decomposition of ${\mathbf{\Xi}}_{\textrm{pst}}^{1}(X_{k})$. This covariance matrix encodes the one-step-ahead forecasting error of $\mathbf{k}_{Z}(X_{k})$ conditioned on $J_{k-1}$. Given the operator-valued kernel defined earlier, in the current setting this matrix is ${\mathbf{\Xi}}_{\textrm{pst}}^{1}(X_{k})={\kappa}_{\textrm{pst}}(X_{k},X_{k})\,\mathbf{K}_{\textrm{bc}}$. Now recall that the expression for the sparse posterior kernel (see (A.6) in Appendix A.3) is
@@
where $\tilde{\mathbf{B}}\coloneqq\left(\mathbf{K}_{ZZ}^{-1}-\sigma_{Y}^{2}\tilde{\mathbf{C}}_{XX}^{-1}\right)$. Hence, the single-step prediction error is represented by the discretized version of the posterior kernel $\xi_{\textrm{pst}}(\cdot,\cdot)$ and subsequently the operator $\mathcal{C}_{\textrm{bc}}$ (see Appendix E for details).
@@
$\begin{split}\mathbf{k}_{Z}(X_{k})&=\mathbf{A}_{\kappa}^{k}\;\mathbf{k}_{Z}(X_{0})+\sum_{j=0}^{k-1}\mathbf{A}_{\kappa}^{j}\;{\mathbf{L}}_{\textrm{pst}}(X_{k-j-1})\bm{\omega}_{k-j}.\end{split}$
@@
$\begin{split}\hat{\bm{x}}_{k}&=\mathbb{E}\left[X_{k}\mid X_{0}=\bm{x}_{0}\right]=\mathbf{A}_{f}^{k}\;\mathbf{k}_{Z}(\bm{x}_{0}),\end{split}$ (2.10)
where the weight matrix is $\mathbf{A}_{f}^{k}\coloneqq\left(\mathbf{W}\,\mathbf{\Lambda}^{k}\,{\mathbf{V}}_{f}\right)^{\top}\in\mathbb{R}^{D\times M}$. In deriving (2.10) we used (2.7), the relationship $X_{k}=(\mathbf{W}{\mathbf{V}}_{f})^{\top}\mathbf{k}_{Z}(X_{k})$, and the fact that a linear transformation of a Gaussian random vector yields another Gaussian random vector (Peebles Jr, 2001). This last property was the key to calculating the Bayesian-consistency matrix $\mathbf{K}_{\textrm{bc}}\in\mathbb{R}^{M\times M}$ and enforcing the structure of the operator–valued prior covariance function to $\xi_{\textrm{pr}}(\bm{x},\bm{x}^{\prime})\!=\!\kappa_{\textrm{pr}}(\bm{x},\bm{x}^{\prime})\,\mathcal{C}_{\textrm{bc}}$.
@@
$\begin{split}\mathbf{K}_{\textrm{pst}}^{k}(\bm{x})+\sigma_{Y}^{2}\mathbf{I}_{D}&=\mathbf{A}_{f}^{0}\left({\mathbf{\Xi}}_{\textrm{pst}}^{k}(\bm{x})+\sigma_{Y}^{2}\mathbf{K}_{\textrm{bc}}\right)\left(\mathbf{A}_{f}^{0}\right)^{\top}.\end{split}$ (2.11)
Given the independence assumption we placed on the components of the flow map, the single-step covariance matrix with respect to the full-state observables is $\mathbf{K}_{\textrm{pst}}^{1}(\bm{x})=\kappa_{\textrm{pst}}(\bm{x},\bm{x})\,\mathbf{I}_{D}$. Hence, it follows that discretizing the operator $\mathcal{C}_{\textrm{bc}}$ leads to
$\begin{split}\mathbf{K}_{\textrm{bc}}&=\left(\left(\mathbf{A}_{f}^{0}\right)^{\top}\mathbf{A}_{f}^{0}\right)^{\dagger}={\mathbf{V}}_{\kappa}^{*}\left(\mathbf{V}_{f}\mathbf{V}_{f}^{*}\right)^{\dagger}{\mathbf{V}}_{\kappa}.\end{split}$
Where the Hermitian adjoint is indicated by , and since the matrix $\mathbf{K}_{\textrm{bc}}$ always has a rank less than or equal to $D$, we employed the Moore–Penrose pseudoinverse .
@@
$\zeta_{k}\coloneqq\mathbf{k}_{Z}(X_{k})-\mathbf{k}_{Z}(\hat{\bm{x}}_{k})=\sum_{j=0}^{k-1}\mathbf{A}_{\kappa}^{j}\;{\mathbf{L}}_{\textrm{pst}}(X_{k-j-1})\bm{\omega}_{k-j},$
@@
$=\mathbb{E}\left[\sum_{j=0}^{k-1}\mathbf{A}_{\kappa}^{j}\;\mathbf{\Xi}_{\textrm{pst}}^{1}\left(X_{k-j-1}\right)\;\left(\mathbf{A}_{\kappa}^{j}\right)^{\top}\mid J_{k}\right].$
As this expression is in general not analytically tractable, similar to Girard et al. (2002a, b) we can approximate it by performing Monte Carlo sampling or, if efficiency is of greater concern than accuracy, a second-order Taylor series expansion about the predicted mean [^2] The unscented transform is also an alternative and very attractive approach (Julier and Uhlmann, 2004).:
@@
which we can interpret as iteratively accumulating the prediction errors arising from random ‘shocks’ at each intermediate step, and where the magnitude of these shocks are dependent on the system’s state.
@@ 2.5 Model Selection
Hyperparameter optimization proceeds as if one were only interested in learning the optimal feature space for the flow map with the VFE method (Titsias, 2009). This is an optimization problem that is solved by maximizing the sum of the VFEs of each of the components of the flow map $\bm{F}=\{f_{i}\}_{i=1}^{D}$. The reasoning is that the theoretical Bayesian-consistent structure of our model also ensures a coherent parameterization of the kernel Koopman and Perron–Frobenius operators.
@@
Thereby, we used the approximate linear dependence (ALD) selection criterion (Engel et al., 2004) as a first pass to determine a subset of candidate locations for the pseudo-inputs from the training inputs. Since the ALD approach is a local ‘whack-a-mole’ active learning scheme, we followed it up by implementing active learning Cohn (Seo et al., 2000): a more globally oriented methodology that approximates a full A-optimal design (Gramacy, 2020). As a last step, only after a predefined number of new dictionary functions had been identified was the VFE hyperparameter tuning algorithm activated, with the greedy results given as a warm start [^4] One could consider including other Subset of Data algorithms, e.g. (Vincent and Bengio, 2002; Dhillon et al., 2004).. This process is interleaved with the adaptation of generative hyperparameters, and repeats until no more pseudo-inputs are added to the basis dictionary or a threshold is passed, e.g., an upper limit on the dictionary size.
@@ 3 Numerical Experiments
With the foundational unification of GP regression and DMD complete, we present the results obtained from a few experiments. In all of the experiments, we used the $\text{Matérn}^{5/2}$ kernel (Matérn, 2013) with automatic relevance determination (ARD; Rasmussen and Williams (2006)):
@@
For all numerical experiments, the data was standardized, and we emphasize that the data collection setup was asymmetric with respect to measurement noise. Since our model is not theoretically equipped to account for sensor noise on the inputs, we restricted the corruption to the dynamically evolved states. This simplification is admissible in surrogate modeling, but unrealistic in the majority of physical data acquisition contexts.
@@ 3.1 Generalization Errors
We compared the multi-step errors of our algorithm (referred to as GP-TCCA) using the Van der Pol oscillator as a benchmark:
@@
where $\alpha=2$ controls the degree of nonlinearity. The dynamical system was simulated by a numerical ODE solver and sampled in time intervals of $\Delta t=50.0\text{ [ms]}$. Figure 1 depicts an example trajectory of the oscillator and its associated forecast.
@@
Refer to captionRefer to caption
@@
We also compared the VFE method to the ‘variational approach to Markov processes’ (VAMP) for hyperparameter optimization (Wu and Noé, 2020). Specifically we maximized the VAMP-$2$ score of the model.
When optimizing the VAMP score, the two DMD algorithms are virtually indistinguishable, suggesting that the Tikhonov regularization has little to no effect. This indicates that the VAMP-$2$ objective does not capture sensor noise as effectively as maximizing the VFE of the flow map. Figure 2 shows that our GP-DMD variant outperforms Exact EDMD under noisy conditions. This aligns with our expectations given the explicit treatment of sensor noise and the additional structure imposed by the Bayesian model. However, the nonlinear growth did prompt a closer examination of our regularization strategy.
3.2 Decoupling the Lifted Observation Model
Refer to captionRefer to caption